OpenStudy (anonymous):

A certain bacteria population is known to triples every 150 minutes. Suppose that there are initially 110 bacteria. What is the size of the population after t hours?

OpenStudy (anonymous):

You know it is an exponential functions when they say 'doubles', 'triples' for some constant amount of time.

OpenStudy (anonymous):

Exponential functions come in the form: \[\Large ce^{kt}\]You have to use the information they give you to solve for \(c\) and \(k\).

OpenStudy (anonymous):

may i have more details.... i thought the answer is fx=110*3^(t/2/5)..... am i right?

OpenStudy (anonymous):

Usually something like 'initially' or 'at the start' is a good indicator for solving for \(c\). Since at that time \(t=0\).

OpenStudy (anonymous):

t/2/5? Is that \(\frac{t}{2/5}\) or \(\frac{t/2}{5}\)?

OpenStudy (anonymous):

the first one, 

OpenStudy (anonymous):

It looks good. Why don't you just check?

OpenStudy (anonymous):

When \(t=0\) does it equal \(110\)?

OpenStudy (anonymous):

Does \(\frac{f(t+2.5)}{ f(t)} = 3 \)?

OpenStudy (anonymous):

These are ways you would check!

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

ok, i am double right now, thank you

OpenStudy (anonymous):

What language do you normally speak?

OpenStudy (anonymous):

yes, When t=0 does, the bacterial should equal 110

OpenStudy (anonymous):

chinese

OpenStudy (anonymous):

why you are asking?

OpenStudy (anonymous):

Maybe someone who speaks Chinese would see and help too. Who knows.

OpenStudy (anonymous):

Now did you try to divide \(\frac{ f(t+2.5) }{ f(t) } \)?

OpenStudy (anonymous):

you are right, the problem is not language, it's i just totally did not remember exponential question anymore

OpenStudy (anonymous):

well,f(t+2.5)f(t) equals 3. so, it's triple from 110 to 330, right?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Let me see...

OpenStudy (anonymous):

may i ask another question about cubic function?

OpenStudy (anonymous):

\[\huge\frac{110\cdot 3^{\frac{t+2.5}{2.5}}}{110\cdot 3^{\frac{t}{2.5}}}\]\[\huge=\frac{110\cdot 3^{\frac{t}{2.5}+1}}{110\cdot 3^{\frac{t}{2.5}}}\]\[\huge=3\cdot \frac{110\cdot 3^{\frac{t}{2.5}}}{110\cdot 3^{\frac{t}{2.5}}}\]\[\huge=3\]So it's correct.

OpenStudy (anonymous):

What about cubic function?

OpenStudy (anonymous):

Find the equation for the cubic function, f(x), with roots at 5, 6 and -5, and has a y-intercept at (0, 13).

OpenStudy (anonymous):

So, remember that \[(x-r_1)(x-r_2)(x-r_3)=0\]Where \(r\) are roots.

OpenStudy (anonymous):

But even after that, we have to compress/expand it so that the y-intercept is correct.

OpenStudy (anonymous):

so, what i got f(x)=x^3-6x^2-25x+13, I try soooo long, still got the same answer, and after I double check, the answer is wrong

OpenStudy (anonymous):

You need to compress it

OpenStudy (anonymous):

yesss. you are right about the root formular

OpenStudy (anonymous):

Now, what is the current y intercept? What happens when \(x=0\)?

OpenStudy (anonymous):

when x=0, y=13?

OpenStudy (anonymous):

\[c(x-5)(x-6)(x-(-5))=0\]Okay so \(c\) is some constant... ok?

OpenStudy (anonymous):

We need to solve for \(c\). It will compress our function. To solve for \(c\), plug in the point \((0,13)\) Like this: \[c((0)-5)((0)-6)((0)-(-5))=13\]

OpenStudy (anonymous):

Now can you solve for \(c\)?

OpenStudy (anonymous):

ok...is c= 13/150 or c=13-150

OpenStudy (anonymous):

c=13/150, right?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

So our function is just: \[y=\frac{13}{150}(x-5)(x-6)(x+5)\]But you need to expand the polynomials.

OpenStudy (anonymous):

yes, sounds great! now, i got it,

OpenStudy (anonymous):

Do you understand how I did that?

OpenStudy (anonymous):

I don't remember much anything about cubic or exponential functions. I did it by using rules of algebra.

OpenStudy (anonymous):

yes, i indeed understood, thanks